# A Survey on Sparse Autoencoders: Interpreting the Internal Mechanisms of Large Language Models

### v1

@@ 3.2.1 Improve Architecture
Gated SAE.   The Gated SAE Rajamanoharan et al. (2024a) is a modification of the standard SAE that aims to improve the trade-off between reconstruction fidelity and sparsity enforcement. Traditional SAEs suffer from shrinkage bias Wright and Sharkey (2024), where the $L_{1}$-norm regularization systematically underestimates feature activations, leading to reduced reconstruction accuracy. Inspired by Gated Linear Units Dauphin et al. (2017); Shazeer (2020), Gated SAEs replace the standard ReLU encoder with a gated ReLU encoder, which allows the L1 penalty to be applied solely to the selection mechanism:
@@
To optimize the Gated SAE, the authors introduce an auxiliary loss $\|\mathbf{z}-\hat{\mathbf{z}}_{\text{frozen}}\left(\text{ReLU}\left(\pi_{\text{gate}}(\mathbf{z})\right)\right)\|_{2}^{2}$ on top of the traditional loss function. This addresses the issue of the non-differentiability of the Heaviside step function used in the gating mechanism, and enables gradient flow during backpropagation while still enforcing sparsity. Here, $\hat{z}_{\text{frozen}}$ is a copy of the decoder with frozen weights.
@@
3.2.2 Improve Training Strategy
@@
$\mathcal{L}_{\text{MFR}}=\alpha\left(\frac{1}{N(N-1)}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}(1-\text{MMCS}(W^{(i)},W^{(j)}))\right),$ (20)
where $W^{(i)}$ and $W^{(j)}$ are the decoder weight matrices of different SAEs. Mean of Max Cosine Similarity (MMCS) measures the degree of alignment between the learned features across SAEs. $\alpha$ is a hyperparameter that controls the strength of the regularization. This mutual feature regularization is then combined with the traditional SAE loss to form the final training objective of the Feature Aligned SAE.
@@ 3.3 SAE Training
4 Explainability Analysis of SAEs
4.1 Input-based Explanations
where $f_{<l}(x)$ indicates generating the hidden representation of input text $x$ at the $l$-th layer, and $l$ is the layer our SAE is trained for. This strategy is reasonable for interpreting weight vectors of SAEs because of the sparse nature of SAEs, which indicates that a learned feature vector should only be activated by a certain pattern/concept. Therefore, summarizing the text spans that could maximally activate a certain weight vector gives us a clue to understanding the semantic meaning of the learned feature vector.
[…]
4.2 Output-based Explanations
VocabProj.   Output-based explanations project the learned feature vectors to the output word embeddings of texts to compute the activations. Mathematically, $f_{\text{out}}(w):\mathcal{V}\rightarrow\mathbb{R}^{d}$ denotes the output word embedding layer that returns the output embeddings of word $w$, and we can collect the natural language explanations by:
This mapping process makes sense for decoder-only LLMs because the layers in such models share the same residual stream, enabling the representations in the intermediate layers to be linear correlated to the output word embeddings Nostalgebraist (2020). Recent works Wu et al. (2025b); Gur-Arieh et al. (2025) find that output-based explanations show a stronger promise in interpreting LLM behaviors (i.e., generated texts) compared to input-based ones.
where $\text{MI}(\cdot;\cdot)$ indicates mutual information between two variablesCover (1999), and $U(\mathcal{C})$ includes all possible vectors that express the knowledge $\mathcal{C}$. Practically, the conditional probabilities can be estimated by:
Compared with VocabProj, that only considers $p(w|\mathbf{w}_{m})$, this mutual-information-driven objective highlights the need to normalize the raw activation with $p(\mathbf{w}_{m}|w)$. That is to say, if a word whose output embedding consistently activates various learned feature vectors, it has no specific to interpret any of them.
5 Evaluation Metrics and Methods
5.1 Structural Metrics
Structural metrics focus on assessing whether an SAE behaves as intended during training. SAEs are designed to optimize both reconstruction fidelity and sparsity, as these properties are explicitly enforced in the training loss. Therefore, natural evaluation metrics assess reconstruction accuracy and sparsity in the model’s learned representations.
Reconstruction Fidelity.   The most fundamental way to evaluate reconstruction fidelity is through Mean Squared Error (MSE) and Cosine Similarity Ng et al. (2011), which directly compare the original activations with SAE-reconstructed activations. Additional metrics such as Fraction of Variance Unexplained (FVU) (also known as normalized loss) Gao et al. (2025) and Explained Variance Karvonen et al. (2024) measure how much variance in the original data is retained after SAE reconstructs. Beyond direct reconstruction comparisons, researchers also evaluate how SAEs affect the probability distribution of model outputs. Cross-Entropy Loss Shannon (1948) and KL Divergence Kullback and Leibler (1951) measure the shift in probability distributions when substituting original model activations with SAE-generated activations. If the SAE faithfully reconstructs activations, the probability distributions should remain similar. Similarly, Delta LM Loss Lieberum et al. (2024) quantifies the difference between the original language model loss and the loss incurred when replacing activations with those from the SAE. Another important aspect of reconstruction fidelity is magnitude preservation. The L2 Ratio Karvonen et al. (2024) compares the Euclidean norms of different activations to ensure that the SAE does not systematically alter activation magnitudes.
5.2 Functional Metrics
6 Applications in Large Language Models
6.1 Model Behavior Anlysis
A recent study utilizes SAEs to reveal the mechanism of hallucination, where entity recognition plays a pivot role in recalling facts. A direction distinguishing whether the model knows an entity is identified, which is usually used for hallucination refusal in chat models Ferrando et al. (2025). Some studies focus on interpreting how in-context learning (ICL) is performed within models. One study focuses on general ICL tasks, and task-related function vectors has been successfully isolated Kharlapenko et al. (2024). Another study focuses on reinforcement learning(RL) tasks. Their experiment shows that LLM’s internal representations are capable of capturing temporal difference errors and Q-values that are essential in RL computations Demircan et al. (2025). Besides, one study attempts to examine the working mechanism of instruction following. Their analysis on translation tasks shows that instructions are composed of multiple relevant concepts rather than individual ones He et al. (2025).
Moreover, SAEs have been employed to study behaviors related to toxicity, sycophancy, refusal, and emotions. One recent study shows that features captured by SAEs can be used to construct probes to classify cross-lingual toxicity Gallifant et al. (2025). By examining SAE latents that activate on anger-related tokens, researchers have identified steering vectors that control angry outputs Nanda et al. (2024). Additionally, other research demonstrates that SAEs can reconstruct vectors responsible for refusing to answering harmful questions as well as directions that produce sycophantic responses neverix et al. (2024).
6.2 Model Steering
Unlike supervised concept vectors, such as probing classifiers Belinkov (2022); Zhao et al. (2025); Jin et al. (2025), SAEs can simultaneously learn a large volume of concept vectors. The learned vectors can then be utilized to steer model behaviors in ways similar to supervised vectors (See Figure 3).
As previously mentioned, SAE latents can be employed to produce steering vectors that control model outputs related to toxicity, sycophancy, refusal, and emotions. Other steering applications have also proven feasible. Recently, a study shows that SAE latents are able to capture instructions such as translations. These identified latents are effective in manipulating models to translate inputs according to specific instructions He et al. (2025). Another investigation focusing on semantic searching demonstrates that SAEs can learn fine-grained semantic concepts at various levels of abstraction. These concept vectors can be used to steer models toward related semantic O’Neill et al. (2024). Alternatively, SAEs trained on biological datasets can provide biology-related features that enable the unlearning of biology-relevant knowledge with fewer side effects than existing techniques Farrell et al. (2024). Given that steering effects are generally challenging to control, SAE-TS further utilizes SAE latents to optimize steering vectors by measuring the changes in SAE feature activation caused by steering, thereby helping construct vectors that target relevant features while minimizing unwanted effects Chalnev et al. (2024). Moreover, explanations based on SAE latents risk prioritizing linguistic features over semantic meaning. Wu et al. (2025b) propose a novel approach that promotes diverse semantic explanations, which has been demonstrated to enhance model safety.
6.3 Model Training
@@
SAEs are trained to obtain more sparse and interpretable features. The learned concepts and sparsity are both beneficial in model transparency, which can be utilized in model training to align model with human understanding and improve model performance. Since SAEs includes feature-level constraints, Yin et al. (2024) leverage these constraints to enable sparsity-enforced alignment in post-training. Their experiments demonstrate that this approach achieves superior performance across benchmark datasets with reduced computational costs. Similarly, combine learned concepts with next-token prediction training to build more transparent models. Specifically, they extract influential concepts on outputs from SAEs, then incorporate these concept vectors into hidden states by modifying token embeddings. Results show that models trained with this method perform better and exhibit greater robustness on token prediction and knowledge distillation across benchmark datasets Tack et al. (2025). Moreover, SAEs’ ability to provide large-scale explanations has been well explored. By examining the diversity of activated features,  Yang et al. (2025) developed a new approach to augment data diversity. Another work uses task-specific features learned in SAEs to mitigate unintended features within models, significantly improving model generalization on real-world tasks Wu et al. (2025a).
7 Research Challenges
@@
7.1 Incomplete Concept Dictionary
@@
7.2 Lack Theoretical Foundations
@@ 7.4 Computational Burden
7.5 Connection to the Broader Field of Interpretability
The field of MI has been critiqued for its insufficient engagement with the broader interpretability and NLP research literature Bereska and Gavves (2024); Saphra and Wiegreffe (2024). Many of the research topics within MI, such as polysemanticity, superposition, and SAEs, have been investigated in prior and concurrent non-MI fields, often under different terminologies while addressing the same fundamental challenges (Saphra and Wiegreffe, 2024; Elhage et al., 2022). For instance, the study of polysemanticity and superposition, which aims to understand how features are encoded in the model activations, have been studied in the context of distributed representations (Hinton, 1984; Mikolov et al., 2013b, a; Arora et al., 2018; Olah, 2023), disentangled representations (Higgins et al., 2018; Kim and Mnih, 2018; Locatello et al., 2019), and concept-based interpretability (Nicolson et al., 2024; Kim et al., 2018). Similarly, SAEs are closely related to and draw inspiration from earlier lines of research on sparse coding and dictionary learning (Olshausen and Field, 1997; Gregor and LeCun, 2010; Faruqui et al., 2015; Subramanian et al., 2018). These methods, like SAEs, posit the feature sparsity hypothesis (Elhage et al., 2022) and aim to learn an overcomplete representation to disentangle the features from activation in superposition. Since these fields pursue similar goals or study the same research problems, the current disconnect causes issues such as missing relevant literature, hindering collaboration, unintentionally redefining established concepts, rediscovering existing techniques, and overlooking well-known baselines. Therefore, it is imperative for the MI community to bridge these gaps and more actively integrate findings from related non-MI research.
8 Conclusions
In this survey, we provided a comprehensive examination of SAEs as a promising approach to interpreting and understanding LLMs. SAEs effectively address the challenge of polysemanticity through learning overcomplete, sparse representations that disentangle superimposed features into more interpretable units. We have systematically explored the foundational principles, technical frameworks, evaluation methodologies, and real-world applications of SAEs in the context of LLM analysis. While SAEs have demonstrated considerable success in revealing the internal mechanisms of LLMs, several challenges remain, including the incompleteness of concept dictionaries, limited theoretical foundations, persistent reconstruction errors, and substantial computational requirements. Despite these challenges, SAEs continue to evolve through architectural innovations and improved training strategies, offering deeper insights into the inner workings of increasingly complex LLMs.

### v2

@@ C.1 Improve Architecture
Gated SAE.   The Gated SAE Rajamanoharan et al. (2024a) is a modification of the standard SAE that aims to improve the trade-off between reconstruction fidelity and sparsity enforcement. Traditional SAEs suffer from shrinkage bias Wright and Sharkey (2024), where the $L_{1}$-norm regularization systematically underestimates feature activations, leading to reduced reconstruction accuracy. Inspired by Gated Linear Units Dauphin et al. (2017); Shazeer (2020), Gated SAEs replace the standard ReLU encoder with a gated ReLU encoder, which separates the roles of detecting which features are active and estimating their magnitudes. Compared to traditional SAEs, this architectural separation addresses the shrinkage effect caused by L1 regularization, which adds a linear penalty to all nonzero activations, leading to systematic underestimation of feature magnitudes.
@@
To optimize the Gated SAE, the authors introduce an auxiliary loss $\|\mathbf{z}-\hat{\mathbf{z}}_{\text{frozen}}\left(\text{ReLU}\left(\pi_{\text{gate}}(\mathbf{z})\right)\right)\|_{2}^{2}$ on top of the traditional loss function. This encourages the gating path to produce useful feature selections for reconstruction, without affecting the learned decoder weights. Here, $\hat{\mathbf{z}}_{\text{frozen}}$ is a copy of the decoder with frozen weights.
@@
C.2 Improve Training Strategy
@@
$\mathcal{L}_{\text{MFR}}=\alpha\left(\frac{1}{N(N-1)}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}(1-\text{MMCS}(\mathbf{W}_{dec}^{(i)},\mathbf{W}_{dec}^{(j)}))\right),$ (25)
where $\mathbf{W}_{dec}^{(i)}$ and $\mathbf{W}_{dec}^{(j)}$ are the decoder weight matrices of different SAEs. Mean of Max Cosine Similarity (MMCS) measures the degree of alignment between the learned features across SAEs. $\alpha$ is a hyperparameter that controls the strength of the regularization. This mutual feature regularization is then combined with the traditional SAE loss to form the final training objective of the Feature Aligned SAE.
@@ C.3 SAE Training
Table 2: Evaluation of SAEs
LLaMa Scope Pythia SAE Gemma Scope
L0:22 L0:41 L0:82 L0:176 L0:445
Structural Sparsity 869.318 112.888 22.141 41.422 80.472 174.74 472.199
MSE 4.9E-5 0.015 2.125 1.836 1.539 1.203 0.707
CE Loss 1.00 0.940 0.974 0.984 0.988 0.993 0.998
KL Div 0.898 -1 0.975 0.984 0.990 0.994 0.997
Variance 0.863 0.918 0.824 0.848 0.875 0.902 0.941
Absorption Mean 2.8E-3 0.227 0.287 0.267 0.105 0.055 0.1347
Full 4.7E-4 0.199 0.333 0.275 0.091 0.038 0.045
SCR Top 5 0.137 0.330 0.206 0.217 0.210 0.184 0.243
Top 50 0.713 0.414 0.376 0.385 0.407 0.417 0.384
Top 500 -0.727 0.232 0.316 0.309 0.359 0.339 0.384
Sparse Probing LLM 0.904 0.922 0.958
SAE 0.885 0.929 0.952 0.955 0.955 0.957 0.958
Appendix D Evaluation and Comparison of SAEs
As shown in Table 2, we evaluate three series of SAEs: LLaMa Scope, Pythia SAE, and Gemma Scope. For LLaMa Scope, we use the llama-scope-lxa-8x SAE trained on layer 12 of the LLaMa 3.1 8B model. For Pythia, we evaluate the sae-bench-pythia70m-sweep-standard-ctx128-0712 SAE, trained on layer 5 of the Pythia 70M model. Within the Gemma Scope, we conduct an internal comparison across five gemma-scope-2b-pt-res SAEs, all trained on layer 12 of the Gemma2 2B model. These five SAEs differ only in their training L0 sparsity settings, meaning they were trained to activate different numbers of latents per input.
We report both structural and functional evaluation metrics. The structural metrics, listed at the top, include L0 Sparsity, Mean Squared Error (MSE), Cross-Entropy Loss, KL Divergence, and Explained Variance. The functional metrics, shown below, including Absorption, Spurious Correlation Removal (SCR), and Sparse Probing. In the Absorption metric, mean absorption captures the fraction of cases where the correct SAE latents fail to activate for a known feature, while another latent with similar semantics activates instead. Full absorption refers to the stricter case where none of the correct latents activate and the feature is entirely absorbed by an unintended latent. In SCR, we assess how well the SAE reduces unwanted correlations. The Top-5, Top-50, and Top-500 SCR scores measure how much debiasing can be achieved by removing top 5, 50, 500 latents respectively. Finally, Sparse Probing evaluates whether the SAE has learned interpretable and disentangled concepts. We compare the performance of sparse probes using SAE activations against a baseline probe using the LLM’s original dense activations. Higher score indicates that the SAE has successfully learned concept-specific features.
Shown in Table 2, the LLaMa Scope SAE shows significantly higher L0 sparsity (869.3) compared to the others. The Gemma Scope SAEs span a range of sparsities, from highly sparse (L0:22 with 22.1 active latents) to relatively dense (L0:445 with 472.2). Within the Gemma Scope series, we observe that as sparsity denser, the MSE also decreases, indicating improved reconstruction accuracy. Similarly, both the cross-entropy loss and KL divergence scores increase toward 1, suggesting that denser SAEs preserve the model’s original predictive behavior more. The explained variance also improves with denser sparsity, rising from 0.824 at L0:22 to 0.941 at L0:445, which shows that denser SAEs are better at capturing the variance of the original activations.
In functional metrics, LLaMa Scope exhibits extremely low mean and full absorption values (0.0028 and 0.00047), likely due to its high redundancy from overactivation. Among the Gemma Scope models, L0:176 achieves the lowest mean (0.055) and full (0.038) absorption, suggesting a good trade-off between sparsity and disentanglement. Later, absorption increases again at L0:445, indicating that overly dense SAEs may begin to recombine multiple concepts into the same latent. Looking at SCR, all SAEs’ performance improves with top-k increase. However, SCR scores begin to decrease at Top-500, indicating that over-ablation may unintentionally remove useful features. For Sparse Probing, we can see that across all models, SAE probes perform nearly as well as LLM-based probes, with accuracies consistently above 0.95 for Gemma Scope. Notably, the Pythia SAE and Gemma Scope L0:445 achieve sparse probing accuracies that even exceed the LLM baseline, suggesting that these models successfully capture clean, task-relevant features using a small set of latents.
Appendix E Other Applications of SAEs in LLMs
E.1 Model Training
@@
SAEs are trained to obtain more sparse and interpretable features. The learned concepts and sparsity are both beneficial in model transparency, which can be utilized in model training to align the model with human understanding and improve model performance. Since SAEs include feature-level constraints, Yin et al. (2024) leverage these constraints to enable sparsity-enforced alignment in post-training. Their experiments demonstrate that this approach achieves superior performance across benchmark datasets with reduced computational costs. Similarly, combine learned concepts with next-token prediction training to build more transparent models. Specifically, they extract influential concepts on outputs from SAEs, then incorporate these concept vectors into hidden states by modifying token embeddings. Results show that models trained with this method perform better and exhibit greater robustness on token prediction and knowledge distillation across benchmark datasets Tack et al. (2025). Moreover, SAEs’ ability to provide large-scale explanations has been well explored. By examining the diversity of activated features,  Yang et al. (2025) developed a new approach to augment data diversity. Another work uses task-specific features learned in SAEs to mitigate unintended features within models, significantly improving model generalization on real-world tasks Wu et al. (2025a).
Appendix F SAE and Probing-Based Methods
In addition to SAEs, probing-based methods have emerged as a prominent family of techniques for interpreting and steering the internal representations of LLMs Giulianelli et al. (2018); Vig et al. (2020); Meng et al. (2022); Guerner et al. (2023); Geiger et al. (2023). These methods operate under the assumption that meaningful concepts are linearly separable within the model’s representation space Mikolov et al. (2013b); Pennington et al. (2014); Park et al. (2023); Nanda et al. (2023). Probing formulates steering as a supervised learning problem, which given a small set of labeled examples, the method learns low-rank projection vectors that can both detect and influence the presence of target concepts during generation.
Based on the recent research, SAEs show certain limitations compared to linear probes for language model interpretation. First, the AXBENCH study Wu et al. (2025c) demonstrates that even simple linear probe baselines like difference-in-means consistently outperform SAEs on concept detection tasks, with SAEs falling behind on both concept detection and model steering benchmarks. Second, Kantamneni et al.’s comprehensive evaluation Kantamneni et al. (2025) across 113 probing datasets reveals that SAEs fail to provide consistent advantages in challenging regimes like data scarcity, class imbalance, label noise, and covariate shift. There remains a long way for the SAE research community to go in making sparse autoencoders more robust, generalizable, and useful in practical applications before they can reliably outperform simpler interpretability approaches.
Appendix G Research Challenges
@@
G.1 Incomplete Concept Dictionary
@@
G.2 Lack Theoretical Foundations